# SGLDBench: A Benchmark Suite for Stress-Guided Lightweight 3D Designs

### v1

@@ 3.2 Stress simulation
At the core of topology optimization and lattice infill methods is the numerical simulation of a stress tensor field using the selected boundary conditions and material properties. SGLDBench provides a MATLAB-interfaced C++ implementation of a multigrid elasticity solver to efficiently simulate the stress field. The implementation uses a geometric multigrid solver as a preconditioner for a conjugate gradient method to solve a sparse linear system of equations, i.e., $KU=F$. The global stiffness matrix $K$ is assembled from the element stiffness matrices under the assumption of a linear material law. The computation of the element stiffness matrices considers the stiffness tensor to reflect the material properties and the strain matrix to express the strain-stress relationship. $U$ is the static displacement vector responding to the external loads $F$. A number of previous works have addressed the efficient assembly of the system matrix $K$ as well as the specific adaptations of numerical solvers for linear elasticity simulation in the context of topology optimization  [30, 31, 32, 33, 37].
SGLDBench’s multigrid implementation follows mostly the one by Wu et al. [31], in that it assembles numerical stencils on-the-fly and performs multigrid interpolation and restriction across multiple levels simultaneously. However, since the implementation is designed for MATLAB running on a CPU, some of the internal data and computation layouts have been changed. Firstly, SGLDBench uses MATLAB’s built-in Cholesky solver instead of the Cholesky solver from the TAUCS library for solving the linear system on the coarsest multigrid level. Secondly, SGLDBench switches from a matrix-free node-based computation layout to a matrix-free element-based one to leverage MATLAB’s powerful matrix-vector operations. Instead of assembling the stencils per grid vertex on-the-fly via indexed memory access operations, SGLDBench builds a generic element matrix and uses MATLAB to efficiently compute the products of this matrix and the 8-node displacement vectors of each single element. Since the stiffness matrix of an element with density $\rho$ is obtained by correspondingly scaling the generic stiffness matrix, the final results only need to be scaled by $\rho$. For each element, the 8 displacement vectors at the vertices are assembled into one column of an element-wise displacement matrix as shown in Fig. 5. MATLAB is then used to compute all matrix-vector products between the generic element matrix and each column in the displacement matrix.
@@ 3.3 Infill computation
For each of the six lightweight design methods provided by SGLDBench, the user selects specific parameters and lets SGLDBench compute the material layout. SGLDBench provides MATLAB code for topology optimization, porous infill optimization, PSL-guided material layout and volumetric Michell trusses. To generate a Voronoi infill, MATLAB calls a Python script including precompiled libraries. Conforming lattice structures are generated by compiling code from a publically available repository and running the executable from MATLAB with the required parameters.
@@ 3.3.1 Topology optimization
Topology optimization minimizes the compliance of a material layout under the constraint of exerting forces and a prescribed global material consumption. SGLDBench implements the density-based topology optimization approach  [34, 38]. To formulate the minimization problem over a discrete set of elements $e$ with pseudo densities $\rho_{e}$, a generated voxel model is used to build a hexahedral finite element discretization of a linear elastic material. The object’s compliance $c$ is computed by summing the strain energy over all material elements, i.e.,
@@
With selected material properties, i.e., Young’s modulus $E_{0}$ of the solid ($\rho_{e}=1$) and the linear material law, topology optimization proceeds in three steps: 1) A large linear system is solved using the MATLAB multigrid implementation to compute the force-induced displacements of the hexahedral vertices. 2) The derivatives of the total strain energy $c$ and the total volume $V$ with respect to the elements’ material variables $\rho_{e}$ are computed and used to guide the material distribution to maximize stiffness. 3) The design is updated according to the computed sensitivities. These steps are repeated until the change in material distribution is below a threshold or the number of iterations reaches the prescribed maximum iterations. The computational pipeline for topology optimization is mainly written in MATLAB, using optimized C++ code for solving the linear system and updating the design variables.
Density-based topology optimization takes the available material budget $V_{0}$ as the constraint, known as the global volume constraint. Thus, the problem-specific constraint function for a material layout $\phi$ using $n$ hexahedral elements is
@@
We use the so-called Solid Isotropic Material with Penalization (SIMP) model, where a non-zero constant minimum value represents the background stiffness of the void region. Topology optimization is conducted in an iterative manner. In each iteration, the design variables are updated within a prescribed step size through a gradient-based optimizer. SGLDBench uses the optimality criteria method to do so [34]. After optimization, the design variables shall converge to a (near-) binary layout that indicates the spatial material distribution.
It’s worth mentioning that several auxiliary processes are also introduced in practical topology optimization for good design quality. For instance, the density-based filtering to counteract numerical instabilities and the Heaviside projection to promote the generation of a binary design, where the proxy density value of each voxel is encoded by the projected value of the filtered value of the design variable [39].
Fig. 6 shows the optimized shapes after different optimization iterations. The optimization produces a mono-scale design comprising mainly a thick resistant strand along the maximum stress directions.
Figure 6: Topology optimization process. (a) to (d): intermediate states of the optimized material field, with per voxel material fill-rate in (0,1). (e) Final binary material field after thresholding.
@@ 3.3.2 Porous infill optimization
Porous infill optimization is an extension of density-based topology optimization. It creates porous substructures that spread across the design domain. This is achieved by replacing the global volume constraint with local volume constraints that prevent materials from accumulating and forming bold solid regions. Specifically, the global volume constraint restricts the total material consumption in the simulation domain. In contrast, the local volume constraint enforces an upper bound $V_{e}$ for the percentage of solid voxels over all voxels in a prescribed neighborhood of voxel $e$. Apart from this, the optimization process follows mostly the one used by density-based topology optimization.
For the material around at each voxel $e$, the local volume constraint leads to the constraint function
@@
$R_{e}$ is the radius of a spherical region centered at a voxel’s center in which the local material accumulation is counted. SGLDBench’s implementation of density-based topology optimization in MATLAB has been extended about the specific constraint function of porous infill optimization. The method of moving asymptotes (MMA) is used as the optimizer to update the design variables [40].
Fig. 7 shows the optimization process of porous infill optimization. In contrast to standard topology optimization, porous infill optimization produces a space-filling multi-scale design. Such designs show lower stiffness than those produced by topology optimization subject to a global volume constraint since some material is deposited in regions that do not contribute significantly to the overall stiffness. On the other hand, in general, such designs are more robust with respect to varying load conditions and local damage [7, 9]. Furthermore, porous infill optimization in 3D tends to form wall-like structures that are along the major and minor principal stress directions, see Fig. 2b.
Figure 7: Porous infill optimization process. (a) to (d): intermediate states of the optimized material field, with per voxel material fill-rate in (0,1). (e) Final binary material field after thresholding.
@@ 3.3.3 Voronoi Infill
where $\hat{r}$ the size of the largest radius (determined as a fraction of the objects bounding box diagonal length) and $\rho\in(0,1]$ gives the ratio of smallest to largest radii. As the von Mises stress has an arbitrary range from smallest to largest values with spatially very concentrated extremes, we normalize and homogenize the field using an inverse cumulative distribution function (icdf).
@@
As seen in Fig. 3, the edges of the Voronoi infill do not follow the direction of stresses in the initial solid object. Whereas the Delaunay criterion guarantees the most regular simplices when applied on the available sampling points, there is no trivial control for edge directions in the Voronoi graph, for instance, to construct Voronoi infills with controlled elasticity  [41] or constraint alignment of the Voronoi edges  [46].
SGLDBench provides Voronoi infill generation via Python due to the easy accessibility of required functionality. The SciPy [47] library is used to generate the Voronoi and Delaunay graphs using Qhull [48] and further provides the kd-tree acceleration structure for the Poisson disk sampling. Our code includes fallback methods required for restricting the graph structure to the object domain. See Fig. 8.
@@ 3.3.4 PSL-guided lattice infill
The most straightforward approach to create an infill that considers the principal stress directions in the loaded solid domain is to deposit material along the principal stress lines (PSLs), i.e., the mutually orthogonal trajectories following the major, minor, and medium eigendirections in the stress tensor field. When using line seeding strategies to obtain an as uniform as possible and domain filling distribution of PSLs  [49, 9], PSL-guided infills show surprisingly good mechanical properties in 2D domains. In 3D domains, however, many PSLs do not significantly contribute to the infill’s stiffness, and PSLs might travel through space over a long distance before they intersect with any other PSL or attach to the boundary (see Fig. 10a).
On the other hand, when performing material deposition by voxelizing lines with a selected thickness, the thicker the PSL trajectory is, the more connections are generated. This increases significantly the mechanical performance, and it often results in infills that show superior mechanical properties. In addition, since the stress field needs to be computed only once in the solid domain, the computational complexity is significantly reduced.
SGLDBench uses the publically available MATLAB backend of 3D-TSV [35] to generate PSLs in a 3D stress tensor field. The implemented method aims at producing a domain-filling and evenly-spaced set of PSLs. The method starts from a set of domain-filling seed points and iteratively creates PSL from these seeds. All remaining seeds in a certain distance to the PSL are removed to control the sparseness of the resulting PSL distribution. The thickness of PSLs is selected by the user, and for a selected thickness the density of seeded PSLs is iteratively increased until the given volume budget is roughly reached. A PSL might enter into a region where the three principal stress directions are not uniquely defined and exchange their orientation, i.e., around so-called degenerate points [15]. Tracing a PSL stops if the resulting directional change exceeds a given limit, and the PSL is removed to avoid wasting material.
@@ 3.3.5 Conforming lattice structure
Conforming lattice structure originates from the geometry-based structural dehomogenization presented by Wu et al [50]. In a general term, structural dehomogenization[51, 52, 53] diverges from the fine-resolution simulation and optimization used by density-based topology optimization and porous infill optimization by adopting a multi-scale strategy:
Homogenization-based topology optimization. During the initial optimization, a coarse-scale representation is tuned to approach the optimal distribution of material across a structure. This structure doesn’t represent the exact material layout but rather provides a set of specifications to guide the optimal material layout.
@@
Homogenization-based topology optimization provides an orthotropic direction field as the optimized specifications. Dehomogenization extracts a conforming lattice structure with edges aligning with the direction field, and aspect ratios or sizes conveying the associated properties of the corresponding directions. The optimal directions are given by the principal stress directions of the homogenization-based structure layout. SGLDBench directly feds the stress field in the solid domain to the dehomogenization stage, where it is used to generate a stress-aligned lattice structure. The cell sizes of the generated structure can either adapt to the local von Mises stresses or remain nearly constant. In Fig. 9, the adaptive lattice structure is compared to the one with a uniform cell size. Both infills are stress-aligned with the same material consumption, yet one can easily see that the adaptive lattice structure gives significantly improved stiffness.
Figure 9: Conforming lattice structure. (a) Non-uniform infill layout according to principal stress amplitudes. (b) Uniform layout.
The conforming lattice structure circumvents the intersection issue of PSL-guided infills by relaxing the requirement of strictly following the principal stress directions. The method builds upon the field-aligned hex-dominant meshing approach by Gao et al. [10], which uses an orthogonal frame field and aligns the edges of a hexahedral mesh with this field. The frame field is replaced with the field of principal stress directions, and the mechanical anisotropy is considered in the field-aligned parameterization process.
Due to the existence of degenerate points and regions in a stress field, however, for all but simple fields, it is impossible to compute a conforming hexahedral mesh. Conforming lattice structure addresses this limitation by pursuing a local smoothing strategy. The key idea is to exploit the rotational symmetry of principal stress directions to generate a smoothed direction field from which a conforming hexahedral mesh can be constructed. From this field, a position field whose gradient aligns with the adjusted stress directions is computed. Finally, by using the smoothed direction field and the position field in combination, the method constructs a graph structure and extracts the final mesh from this structure. Conforming lattice structure aim at generating a regular mesh, at the same time maintaining the divergence/convergence properties of the underlying stress field. This is achieved by introducing irregular vertex connections in the final structure, resulting in an edge-graph structure.
@@ 3.3.6 Volumetric Michell trusses
Frame Field Smoothing. The algorithm first applies an FEA to compute the stress tensor field, followed by a frame field generation aligned with the principal stress directions. However, to achieve a global parametric structure, this step smooths out tensor field singularities, sacrificing local alignment near degenerate points for overall global consistency. To address this, the method uses Loubignac iterations [54] to smooth out the discontinuous stress field. This iterative method adjusts the stress field to ensure that it becomes continuous across element boundaries, thereby allowing for smoother and more uniform alignment in subsequent steps. Therefore, a smoothness energy function that penalizes sharp changes of frame directions is minimized. This optimization enables a globally smooth frame field that approximates the original stress directions.
@@
Due to the applied smoothing of the initial stress field, volumetric Michell trusses produce a regular hexahedral lattice structure with improved continuity of load transmissions. Fig. 10 compares an initial stress field to the smoothed frame field that is used to generate an infill with this method. The smoothing, on the other hand, can significantly modify the initial stress field, so that the resulting stiffness of a design can also deviate significantly from the optimum.
Figure 10: Volumetric Michell trusses. (a) PSLs in the loaded solid domain. (b) PSLs in the smoothed stress field. (c) Resulting design.

### v2

@@ 3.2 Stress simulation
At the core of TO and lattice infill methods is the numerical simulation of a stress tensor field using the selected boundary conditions and material properties. SGLDBench provides a MATLAB-interfaced C++ implementation of a multigrid elasticity solver to efficiently simulate the stress field.
The implementation employs a geometric multigrid solver as a preconditioner for a conjugate gradient method to solve a sparse linear system of equations, i.e., $KU=F$. The global stiffness matrix $K$ is assembled from the element stiffness matrices under the assumption of a linear material law. The computation of the element stiffness matrices accounts for the stiffness tensor, which reflects material properties, and the strain matrix, which expresses the strain-stress relationship. $U$ represents the static displacement vector in response to the external loads $F$. Several prior works have addressed the efficient assembly of the system matrix $K$ and the specific adaptations of numerical solvers for linear elasticity simulations in TO  [26, 27, 28, 29, 38].
SGLDBench’s multigrid implementation is primarily based on the work of Wu et al. [27], utilizing on-the-fly numerical stencil assembly and multigrid interpolation and restriction across multiple levels simultaneously. However, the implementation has been adapted for MATLAB running on a CPU, resulting in changes to the internal data and computation layouts.
Firstly, SGLDBench uses MATLAB’s built-in Cholesky solver rather than the TAUCS library’s Cholesky solver for solving the linear system on the coarsest multigrid level. Secondly, it transitions from a matrix-free node-based computation layout to a matrix-free element-based layout to take advantage of MATLAB’s efficient matrix-vector operations. Instead of assembling stencils per grid vertex on-the-fly using indexed memory access operations, SGLDBench constructs a generic element matrix and utilizes MATLAB to compute the products of this matrix with the 8-node displacement vectors of each element. Since the stiffness matrix of an element with density $\rho$ is obtained by correspondingly scaling the generic stiffness matrix, the final results only need to be scaled accordingly. For each element, the 8 displacement vectors at the vertices are organized into columns of an element-wise displacement matrix, as illustrated in Fig. 4. MATLAB then computes all matrix-vector products between the generic element matrix and each column in the displacement matrix efficiently.
@@
Figure 5: Material layout optimization with density-based TO (a) and porous infill optimization (b). Compliance history is depicted by the green curves, with violet squares indicating shown states.
@@ 3.3 Infill computation
For each of the six lightweight design methods provided by SGLDBench, the user selects specific parameters and lets SGLDBench compute the material layout. SGLDBench provides MATLAB code for density-based TO, porous infill optimization, PSL-guided material layout and volumetric Michell trusses. To generate a Voronoi infill, MATLAB calls a Python script including precompiled libraries. Conforming lattice structures are generated by compiling code from a publically available repository and running the executable from MATLAB with the required parameters.
@@ 3.3.1 Topology optimization
TO minimizes the compliance of a material layout under the constraint of applied forces and a prescribed global material consumption. SGLDBench implements the density-based TO approach  [30, 39]. To formulate the minimization problem over a discrete set of elements $e$ with densities $\rho_{e}$, a hexahedral finite element discretization of a linear elastic solid material is generated from the voxelized geometry. The object’s compliance $c$ is computed by summing the strain energy over all material elements, i.e.,
@@
With selected measure of a material’s ability to deform under an applied stress, i.e., the Young’s modulus $E_{0}$ of the solid ($\rho_{e}=1$), and the linear material law, TO proceeds in three steps: 1) A large linear system is solved using the MATLAB multigrid implementation to compute the force-induced displacements of the hexahedral vertices. 2) The derivatives of the total strain energy $c$ and the total volume $V$ with respect to the elements’ densities $\rho_{e}$ are computed and used to guide the material distribution to maximize stiffness. 3) The design is updated according to the computed sensitivities. These steps are repeated until the change in material distribution is below a threshold or the number of iterations reaches the prescribed maximum iterations. The computational pipeline for density-based TO is mainly written in MATLAB, using our optimized C++ code for solving the linear system and updating the design variables.
Density-based TO takes the available material budget $V_{0}$ as the constraint, known as the global volume constraint. Thus, the problem-specific constraint function for a material layout $\phi$ using $n$ hexahedral elements is
@@
We use the so-called Solid Isotropic Material with Penalization (SIMP) model, where a non-zero constant minimum value represents the background stiffness of the void region. In each optimization iteration, the design variables are updated within a prescribed step size through a gradient-based optimizer using the optimality criteria method  [30]. After optimization, the design variables shall converge to a (near-)binary layout that indicates the spatial material distribution.
It’s worth mentioning that several auxiliary processes are also introduced in practical TO for good design quality. For instance, the density-based filtering to counteract numerical instabilities and the Heaviside projection to promote the generation of a binary design, where the proxy density value of each voxel is encoded by the projected value of the filtered value of the design variable [40].
Fig. 5a shows the optimized shapes after different optimization iterations. The optimization produces a mono-scale design comprising mainly a thick resistant strand along the maximum stress directions.
@@ 3.3.2 Porous infill optimization
Porous infill optimization is an extension of density-based TO that generates porous substructures distributed across the design domain. This is achieved by replacing the global volume constraint with local volume constraints, which prevent material from accumulating and forming dense, solid regions.
The global volume constraint restricts the total material consumption within the entire simulation domain. The local volume constraint imposes an upper bound $V_{e}$ on the percentage of solid voxels within a prescribed neighborhood of voxel $e$. These local constraints ensure a more evenly distributed material layout, promoting porosity and lightweight design. Beyond this adjustment, the optimization process largely follows the approach used in density-based TO.
For the material around each voxel $e$, the local volume constraint leads to the constraint function
@@
$R_{e}$ is the radius of a spherical region centered at a voxel’s center, It defines the area within which local material accumulation is measured. SGLDBench’s implementation of density-based TO in MATLAB has been extended to include the specific constraint function for porous infill optimization. The optimization process uses the Method of Moving Asymptotes (MMA)  [41] as the optimizer to iteratively update the design variables.
Fig. 5b illustrates the optimization process of porous infill optimization. Unlike standard TO, porous infill optimization generates a space-filling, multi-scale design. These designs generally exhibit lower stiffness compared to those produced by TO with a global volume constraint, as some material is deposited in regions that do not significantly contribute to overall stiffness. However, such designs are typically more robust under varying load conditions and localized damage [31, 33]. Additionally, porous infill optimization in 3D tends to form wall-like structures aligned with the major and minor principal stress directions, as shown in Fig. 1b.
@@ 3.3.3 Voronoi Infill
where $\hat{r}$ the size of the largest radius (determined as a fraction of the objects bounding box diagonal length) and $\rho\in(0,1]$ gives the ratio of smallest to largest radii. As the von Mises stress has an arbitrary range from smallest to largest values with spatially varying concentrated extremes, we normalize and homogenize the field using an inverse cumulative distribution function (icdf).
@@
The edges of a Voronoi infill do not follow the stress directions in the initial solid object. Whereas the Delaunay criterion guarantees the most regular simplices when applied on the available sampling points, there is no trivial control for edge directions in the Voronoi graph, for instance, to construct Voronoi infills with controlled elasticity  [42] or constraint alignment of the Voronoi edges  [47].
SGLDBench provides Voronoi infill generation via Python due to the easy accessibility of required functionality. The SciPy [48] library is used to generate the Voronoi and Delaunay graphs using Qhull [49] and further provides the kd-tree acceleration structure for the Poisson disk sampling. Our code includes fallback methods required for restricting the graph structure to the object domain. The pipeline is illustrated in Fig. 6.
@@ 3.3.4 PSL-guided lattice infill
The most straightforward approach to create an infill that considers the principal stress directions in the loaded solid domain is to deposit material along the PSLs. When using line seeding strategies to obtain an as uniform as possible and domain filling distribution of PSLs  [50, 33], PSL-guided infills show surprisingly good mechanical properties in 2D domains [51]. In 3D domains, however, many PSLs do not significantly contribute to the infill’s stiffness, and PSLs might travel through space over a long distance before they intersect with any other PSL or attach to the boundary (see Fig. 7a).
On the other hand, when depositing material by voxelizing lines with a selected thickness, the thicker the PSLs are, the more connections are generated. This increases significantly the mechanical performance, and it often results in infills that show superior mechanical properties. In addition, since the stress field needs to be computed only once in the solid domain, the computational complexity is significantly reduced.
SGLDBench uses the publically available MATLAB backend of 3D-TSV [36] to generate PSLs in a 3D stress tensor field. It generates a domain-filling and evenly-spaced set of PSLs. The method starts from a set of domain-filling seed points and iteratively creates PSL from these seeds. All remaining seeds in a certain distance to the PSL are removed to control the sparseness of the resulting PSL distribution. The thickness of PSLs is selected by the user, and for a selected thickness the density of seeded PSLs is iteratively increased until the given volume budget is roughly reached. A PSL might enter into a region where the three principal stress directions are not uniquely defined and exchange their orientation, i.e., around so-called degenerate points [10]. Tracing a PSL stops if the resulting directional change exceeds a given limit, and the PSL is removed to avoid wasting material.
@@ 3.3.5 Conforming lattice structure
Conforming lattice structure originates from the geometry-based structural dehomogenization presented by Wu et al [6]. Structural dehomogenization[52, 53, 54] diverges from the fine-resolution simulation and optimization used by density-based TO and porous infill optimization by adopting a multi-scale strategy:
Homogenization-based TO. During the initial optimization, a coarse-scale representation is tuned to approach the optimal distribution of material across a structure. This structure doesn’t represent the exact material layout but rather provides a set of specifications to guide the optimal material layout.
@@
Homogenization-based TO provides an orthotropic direction field as the optimized specifications. Dehomogenization extracts a conforming lattice structure with edges aligning with the direction field, and aspect ratios or sizes conveying the associated properties of the corresponding directions. The optimal directions are given by the principal stress directions of the homogenization-based structure layout.
SGLDBench directly feds the stress field in the solid domain to the dehomogenization stage, where it is used to generate a stress-aligned lattice structure. Figs. 7a, b show the initial stress directions in the solid domain and the conforming lattice structure where the cell sizes have been further adapted to the local von Mises stresses.
The conforming lattice structure addresses the intersection issue found in PSL-guided infills by relaxing the requirement to strictly follow the principal stress directions. This approach builds upon the field-aligned hex-dominant meshing method by Gao et al. [34], which employs an orthogonal frame field to align the edges of a hexahedral mesh with this field. In the conforming lattice structure, the frame field is replaced with the field of principal stress directions, and mechanical anisotropy is incorporated into the field-aligned parameterization process.
Due to the presence of degenerate points and regions in a stress field, it is generally impossible to compute a conforming hexahedral mesh for all but the simplest fields. The conforming lattice structure addresses this limitation by employing a local smoothing strategy. The key idea is to leverage the rotational symmetry of principal stress directions to generate a smoothed direction field, which enables the construction of a conforming hexahedral mesh.
Figure 7: Major (brown) and minor (green) PSLs in the solid object under load (a) and in the smoothed stress field (c), used to generate the designs in (b) and (d), respectively.
From this smoothed direction field, a position field is computed, ensuring that its gradient aligns with the adjusted stress directions. By combining the smoothed direction field and the position field, the method constructs a graph structure and subsequently extracts the final mesh. To preserve the divergence and convergence properties of the underlying stress field, irregular vertex connections are introduced in the final structure, resulting in an edge-graph representation.
@@ 3.3.6 Volumetric Michell trusses
Frame Field Smoothing. The algorithm first applies an FEA to compute the stress tensor field, followed by a frame field generation aligned with the principal stress directions. However, to achieve a global parametric structure, this step smooths out tensor field singularities, sacrificing local alignment near degenerate points for overall global consistency. To address this, the method uses Loubignac iterations [55] to smooth out the discontinuous stress field. This iterative method adjusts the stress field to ensure that it becomes continuous across element boundaries, thereby allowing for smoother and more uniform alignment in subsequent steps. Therefore, a smoothness energy function that penalizes sharp changes in frame directions is minimized. This optimization enables a globally smooth frame field that approximates the original stress directions.
@@
Due to the applied smoothing of the initial stress field, volumetric Michell trusses produce a regular hexahedral lattice structure with improved continuity of load transmissions, as demonstrated in Fig. 7c, d. However, the smoothing process can substantially alter the initial stress field, leading to significant deviations in the resulting design’s stiffness from the optimal value.